OpenStudy (anonymous):

x^2-y^2=2011 x^2-y^2=2011 @Mathematics

hero (hero):

(x+y)(x-y) = 2011

OpenStudy (anonymous):

Proceed :) please :) sorry

hero (hero):

This looks like a circle problem

OpenStudy (anonymous):

Its part of a problem from this txml competition

OpenStudy (anonymous):

what is the question?

hero (hero):

Oh, it's a hyperbola

OpenStudy (anonymous):

if you are supposed to find the integer solutions, then notice that 2011 is a prime number, so the only way that \[(x+y)(x-y)=2011\] is if \[x+y=2011,x-y=1\] and you can then solve for x and y

OpenStudy (anonymous):

it went something like the sum of integers equals the total distance traveled by plane. The difference between the squares of the two integers = 2011

OpenStudy (anonymous):

ok then you are looking for two integers that are one apart and add to 2011 you get 1005 and 1006

OpenStudy (anonymous):

or i guess i should say you get 1006 for x and 1005 for y

hero (hero):

How on earth did you find that so fast?

OpenStudy (anonymous):

How did you find those two numbers?!!

hero (hero):

Oh nevermind. I didn't see that 2011 is a prime number

OpenStudy (anonymous):

because this is sort of the same as asking for a pythagorean triple where one side is 11 you put \[(x+y)(x-y)=11\] meaning \[x+y=11,x-y=1\] and pretty much solve in your head to get \[x=6,y=5\]

OpenStudy (anonymous):

hold on finding the numbers is not the hard part. two numbers that are one apart and add to 2011

OpenStudy (anonymous):

the trick is recognizing that that is what you have to do

hero (hero):

And this the procedure when a number such as 2011 is a prime number, right?

hero (hero):

So if x^2-y^2 = 7, then x = 4 and y = 3?

OpenStudy (anonymous):

right

hero (hero):

:D

OpenStudy (anonymous):

then take \[2\times 4\times 3=24,4^2+3^2=25\] and your right triangle is \[7,24,25\]

OpenStudy (anonymous):

Do you guys take the txml?

hero (hero):

The what?

OpenStudy (anonymous):

i took a taxi once

hero (hero):

lol

OpenStudy (anonymous):

hahha nevermind

hero (hero):

Is txml the new .xml format for excel?

OpenStudy (anonymous):

Nah its a math competition by the math league

hero (hero):

You're supposed to be at the math league man. What are you doing here? You're not going to win

OpenStudy (anonymous):

hahaha no I'm in this after school club and we take these

OpenStudy (anonymous):

How did you guys get so good at math? I love it, but I can't get good at it. I suck at problem solving